OpenStudy (keekee14):

What is the equation of a circle with the center (0,5) and diameter of 14

OpenStudy (anonymous):

diameter = 14 then radius = 14/2 = 7 Now equation of a circle with coordinates of center (h,k) and radius = r is given by - (x-h)^2 + (y-k)^2 = r^2 h = 0 and k = 5. r = 7 Equation = (x-0)^2 + (y-5)^2 = 7^2 => x^2 + (y-5)^2 = 49.